1. Two heating units A and B are connected in parallel across 120 volts supply. Unit B takes twice as much power as A and the total power is 2160 watts. Determine the resistance of A and B in ohms. 
1
Expert's answer
2022-04-15T05:19:01-0400
"P=\frac{V^2}{R}"

"R=\frac{V^2}{P}=\frac{120^2}{2160}=6.67\:\Omega"